OpenStudy (anonymous):

-8(-6) - (-7)(9) + 42/(-7) please help

OpenStudy (anonymous):

Well do the multiplication first

OpenStudy (anonymous):

Is this order of operations?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Well do PEDMAS Parentheses Exponents Multiplication Subtacrtion (left to right) Addition (Left to right)

OpenStudy (anonymous):

I have tried doing the problem on my own and they t urn out to be the wrong answers

OpenStudy (anonymous):

ok well first off what is -8(-6)?

OpenStudy (anonymous):

hint: - times a minus equals a positive

OpenStudy (anonymous):

a positive 48?

OpenStudy (anonymous):

and then -7(9) is -63... so then what would be next?

OpenStudy (anonymous):

Good job!

OpenStudy (anonymous):

so then I put 48-(-63)= 111 then 111+42= 153 so then the next step is where I got stuck!!

OpenStudy (anonymous):

Yes, good job. Alright the next step is hold on and let me look ;)

OpenStudy (anonymous):

48-(-63) is equal to 48+63 because 5--5= 5 plus ten, five minus a minus five equals a postive ten

OpenStudy (anonymous):

Alright then you have 111+ 42/-7 (which is...)

OpenStudy (anonymous):

Solve that. Then add up the total foryour answer.

OpenStudy (anonymous):

Does that make sense? :)

OpenStudy (anonymous):

-8(-6) - (-7)(9) + 42/(-7) I did the multiplication first, then the rest i got 105

OpenStudy (anonymous):

kit kat that's what I didn't understand... is that suppose to be division or a fraction?

OpenStudy (anonymous):

the 42/ (-7) ??

OpenStudy (anonymous):

Katheryn that is 42 divided by -7 which becomes -6

OpenStudy (anonymous):

i got 105

OpenStudy (xxdragoonxx):

(-7)*(9) evaluates to -63 -8*(-6)-(-7)*(9) evaluates to 111 Because of the minus sign 7 becomes - 7 The answer is -7 42/(-7) evaluates to -6 -8*(-6)-(-7)*(9)+42/(-7) Final answer 105